FACTOR COMPLETELY X^3=27
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

X^3=27
x^3-27=0
(x-3)(x^2+3x+9) ans.

Answer by Alan3354(69443)

Equations are solved, polynomials are factored.
There is, "solve by factoring", but that's not just factor.